OpenStudy (anonymous):

square root -3 how do I write it in a+bi format?

OpenStudy (anonymous):

\[\sqrt{-3}\]

OpenStudy (anonymous):

\[i \sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{-3} = \sqrt{-1}*\sqrt{3}=i \sqrt{3}\]

OpenStudy (radar):

Since there is no real component a=0, or if you wish \[0+i \sqrt{3}\]

OpenStudy (anonymous):

0+ i\[\sqrt{3}\]

OpenStudy (radar):

At some point you must realize that\[i ^{2}=-1\]

OpenStudy (anonymous):

Yep the correct is 0+i3√